5% of 612=, 3% of 704=, 5% of 294= , 4% of 631=


1
Expert's answer
2020-02-14T10:56:44-0500

1% = 1/100

So:

5% of 612 = 612*5*1/100=30.6


3% of 704 = 704*3*1/100=21.12


5% of 294 = 294*5*1/100 = 14.7


4% of 631 = 631*4*1/100 = 25.24